User’s Guide to the Event List package^{∗}

Frank G. Bennett, Jr. December 9, 1998

Abstract

This package makes a cronological list of events.

Please see the file calguide.tex for details on the use of this package.

∗This file is version number 2.2. It was last revised on 1997/11/07 10:44:05.